OpenStudy (cookie_monster):

If I am trying to find the width of a football field, and I already know the length and perimeter, what expression would help me find the width?

jigglypuff314 (jigglypuff314):

perimeter of rectangle = 2(length) + 2(width) P=2L + 2W 2W= P-2L W= (P-2L)/2